OpenStudy (anonymous):

simplify the expression (6-√ 11)(6+√ 11)

OpenStudy (kropot72):

Just multiply the two bracketted terms together. This will remove the radical signs and simplify.

OpenStudy (kropot72):

What didyou get when you multiplied the bracketted terms?

OpenStudy (kropot72):

To multiply using the FOIL method: \[F\ pair=6\times 6=36\] \[O\ pair=6\times \sqrt{11}=6\sqrt{11}\] \[I\ pair=6\times -\sqrt{11}=-6\sqrt{11}\] \[L\ pair=-\sqrt{11}\times \sqrt{11}=-11\] When you add these results you get \[36+6\sqrt{11}-6\sqrt{11}-11=you\ can\ calculate\]

jimthompson5910 (jim_thompson5910):

Alternative Way: Use the difference of squares rule \[\Large (a-b)(a+b) = a^2 - b^2\] \[\Large (6-\sqrt{11})(6+\sqrt{11}) = 6^2 - \left(\sqrt{11}\right)^2\] \[\Large (6-\sqrt{11})(6+\sqrt{11}) = 36 - 11\] \[\Large (6-\sqrt{11})(6+\sqrt{11}) = 25\]

jimthompson5910 (jim_thompson5910):

both lead to the same answer, just two different ways to get there

jimthompson5910 (jim_thompson5910):

is it \[\Large w = \sqrt{7w}\] or is it \[\Large w = \sqrt{7}w\] the difference between the two is the placement of w

jimthompson5910 (jim_thompson5910):

is the second w in the square root?

jimthompson5910 (jim_thompson5910):

ok you would square both sides to get \[\Large w^2 = 7w\] what's next?

jimthompson5910 (jim_thompson5910):

\[\Large w^2 = 7w\] \[\Large w^2 - 7w = 0\] \[\Large w(w - 7) = 0\] how about now?

jimthompson5910 (jim_thompson5910):

well let's check

jimthompson5910 (jim_thompson5910):

plug in w = 0 (first possible solution) \[\Large w = \sqrt{7w}\] \[\Large 0 = \sqrt{7*0}\] \[\Large 0 = \sqrt{0}\] \[\Large 0 = 0\] Since that last equation is true, w = 0 is indeed a solution

jimthompson5910 (jim_thompson5910):

I'll let you check w = 7

jimthompson5910 (jim_thompson5910):

you are correct plug in w = 7 (second possible solution) \[\Large w = \sqrt{7w}\] \[\Large 7 = \sqrt{7*7}\] \[\Large 7 = \sqrt{49}\] \[\Large 7 = 7\] that's true, so w = 7 is also a true solution

jimthompson5910 (jim_thompson5910):

you're welcome